OpenStudy (anonymous):

Now it is your turn to plan the trajectories required to launch a spacecraft through a specific route in space. The launch area is identified on the map below. A coordinate plane is shown with a point at 1, 2 labeled, Launch Area. Launch Area:___(1, 2)___ Point A:___(0, 3)___ Point B:___(-3, 0)___ Point C:___(-1, -4)___ 1.Graph the cordinates of the specifics points in space your spacecraft will travel to.

OpenStudy (anonymous):

Launch Area:___(1, 2)___ Point A:___(0, 3)___ Point B:___(-3, 0)___ Point C:___(-1, -4)___ 1.Graph the cordinates of the specifics points in space your spacecraft will travel to. You must show your work on each question below. 2.Determine the equation of the line, in standard form, that will get your spacecraft from the Launch Area to Point A. 3.Determine the equation of the line, in point-slope form, that will get your spacecraft from Point A to Point B. 4.Determine the equation of the line, in slope-intercept form, that will get your spacecraft from Point B to Point C. 5.Convert the equations you arrived at in question 2 and into slope-intercept form. Make sure to include all of your work. 6.Reflect back on this scenario and each equation you created. Would any restrictions apply to the domain and range of those equations? Explain your reasoning using complete sentences.

OpenStudy (anonymous):

I got u

OpenStudy (anonymous):

Thanks, do you understand it?

OpenStudy (anonymous):

I got it

OpenStudy (anonymous):

i think so

OpenStudy (anonymous):

But you go step by step with me? I want to be able to understand it

OpenStudy (anonymous):

yea i am reading it

OpenStudy (anonymous):

So? @Tyrone56

OpenStudy (anonymous):

@sierraleone17

OpenStudy (anonymous):

standard form is Ax+By=C slope intercet is y=mx+b point slope is y1-y2=x2-x1 not sure on point slope... after u read these equations re-read the determine statements. hope it helps

OpenStudy (anonymous):

lol idk

OpenStudy (anonymous):

you guys help help me out step by step?

OpenStudy (anonymous):

i just gave you something to go by